Merk has 580 million shares outstanding that currently trade at $374.

Part 1

If Merk pays a 70% stock dividend, what will be the new share price?
Answer #1

Number of shares outstanding after the stock dividend = (Number of shares before the stock dividend) * (1 + Stock dividend %)
Number of shares outstanding after the stock dividend = (580*1000000)*(1+70%) = 986000000

Market capitalization (or the market value of a company) = (Number of shares outstanding)*(Share price)
Now, market value of a company does not change due to stock dividend.
(Share price before stock dividend)*(Number of shares before stock dividend) = (Share price post dividend)*(Number of shares post stock dividend)
Substituting the values, we get;
374*(580*1000000) = (Share price post dividend)*(986000000)
374*(580*1000000)/(986000000) = Share price post dividend
=>Share price post dividend=374*(580*1000000)/(986000000)
=374*(580)/(986)
=216920/(986)
=220

Answer: New share price=$220